tjjohnson2001:

Solve for x: 2 over 5 plus 3 over 5x equals x plus 5 over 10? x = -3 x = 2 x = -3 and x = 2 No Solution

Narad:

The equation is \[(2/5)+(3/5x)=(x+5)/10\] \[(2x+3)/(5x)=(x+5)/10\] \[2x+3)/x=(x+5)/2\] \[2(2x+3)=x(x+5)\] \[4x+6=x^2+5x\] \[x^2+x-6=0\] Factorising \[(x-2)(x+3)=0\] Finally, \[x=2\] and \[x=-3\] The answer is x=2 and x=-3